OpenStudy (anonymous):

How do I find the missing values that solve this equation. y = 3x (0, ____) (-1,____) (7,____)

mathslover (mathslover):

ok so if y = 0 then 0 = 3x

mathslover (mathslover):

hence for y = 0 , x = 0 right ?

OpenStudy (anonymous):

Yes.

mathslover (mathslover):

ok so if y = -1 can u tell me what will be x ? -1 = 3x ... x =?

OpenStudy (anonymous):

would it be 1?

mathslover (mathslover):

divide both sides by 3 .. \[\frac{-1}{3}=\frac{3x}{3}\] \[\frac{-1}{3}=\frac{\cancel{3}x}{\cancel{3}}\]

OpenStudy (anonymous):

That gives me 3 right?